# Decide less, communicate more: On the construct validity of end-to-end fact-checking in medicine

### v2

@@ Abstract
Technological progress has led to concrete advancements in tasks that were regarded as challenging, such as automatic fact-checking. Interest in adopting these systems for public health and medicine has grown due to the high-stakes nature of medical decisions and challenges in critically appraising a vast and diverse medical literature. Evidence-based medicine connects to every individual, and yet the nature of it is highly technical, rendering the medical literacy of majority users inadequate to sufficiently navigate the domain. Such problems with medical communication ripens the ground for end-to-end fact-checking agents: check a claim against current medical literature and return with an evidence-backed verdict. And yet, such systems remain largely unused. To understand this, we present the first study examining how clinical experts verify real claims from social media by synthesizing medical evidence. In searching for this upper-bound, we reveal fundamental challenges in end-to-end fact-checking when applied to medicine: Difficulties connecting claims in the wild to scientific evidence in the form of clinical trials; ambiguities in underspecified claims mixed with mismatched intentions; and inherently subjective veracity labels. We argue that fact-checking should be approached and evaluated as an interactive communication problem, rather than an end-to-end process. Our data and code is available at https://github.com/SebaJoe/decide-less-communicate-more.
@@ 1 Introduction
Decision making in medicine is personal, intimate, and high-stakes. Traditionally the patient—often a lay person unfamiliar with medicine—converses with their care providers about questions about their health. However, the reality is far from this picture: most Americans resort to the web when they have a health-related question (Fox and Duggan, 2013). Today, social media and AI have made medical knowledge seemingly accessible. But claims made by others on the web (or by an chatbot) can be inaccurate or inapplicable. This, combined with eroding health literacy (Champlin et al., 2017), have led to challenges in public health (Hassan et al., 2015) as well as patient-provider sessions.
@@
To assess the boundary of AI-driven fact-checking systems, we focus on real-world medical claims from social media, preserving their original context. We contend that fact-checking systems should mirror how experts (e.g., physicians, care providers) evaluate and respond to such claims. We designed an annotation study that examines how medical experts verify claims using retrieved medical evidence. Experts were asked to assess medical claims present on social media, i.e., Reddit forums about a particular medical condition) by synthesizing retrieved randomized controlled trial (RCT) abstracts and explaining their judgments. This provides a novel benchmark for systems verifying “in the wild” claims. However, we highlight fundamental obstacles that challenge the construct validity of end-to-end automated systems for fact-checking: given a claim, provide a veracity judgment. We identify inherent difficulties in this setup for even domain experts, including: connecting claims with evidence; ambiguity from underspecified claims leading to valid yet contradictory interpretations; and challenges in achieving annotator consensus due to the inherent subjectivity of veracity labels. These issues suggest that the existing framing fact-checking of as an end-to-end classification task is inadequate for real-world settings, which may explain in part why such systems have not been put into wide use.
2 Background: medical claim-checking
Dataset Domains Source Labels Expl. Evidence Type Claim Example
@@ 2 Background: medical claim-checking
Table 2: A comparative view of related work in medical/health fact-checking. N/A indicates that the component is not applicable to the task (e.g., RedHOT does not perform claim verification).
@@ 3 Case Study Methods
Figure 1: An overview of our AI-in-the-loop expert study pipeline: given a claim from a subreddit, we extract the PIO elements and retrieve the evidence automatically. The evidence, its context, and the evidence are then presented to a medical expert to provide a judgment and a rationale for the factuality of the claim.
@@
We formulate the following study for expert claim verification, in accordance to the task construction for automatic fact-checking systems discussed in section 2. We use claims from the Reddit Health Online Talk (RedHOT) corpus (Wadhwa et al., 2023), which contains 22,000 annotated posts from Reddit across 24 health conditions. RedHOT defines a claim as a statement indicating (often only implicitly) a causal relationship between an intervention and an outcome. To help experts contextualize these claims, we provide the full post and Population, Intervention, Outcome (PIO) descriptors (illustrated in Table 3. See Appendix D,Appendix E for derivation details.) Note that we leave off the Comparators because in practice claims rarely mention the comparator explicitly (e.g., “Vitamin C cured my flu”). Given a contextualized medical claim, a medical expert constructs a hierarchy of evidence (Guyatt et al., 2008a, b, 2011) based on relevance and quality, assesses the claim’s veracity, and provides a grounded explanation.
@@
The overall task, without any aid, places a high cognitive load on experts who must comb through and synthesize multiple pieces of evidence (Juneja and Mitra, 2022). With the aim of easing this, we automated several steps of the fact-checking process and integrated them into an intuitive web-based annotation interface (see Appendix Q). These features enable experts to focus on critical aspects: evaluating evidence relevance and synthesizing it to support or refute claims. Detailed annotation guidelines are provided in Appendix G.
@@
For each claim, we used its PIO elements to automatically retrieve ten published relevant RCT abstracts from Trialstreamer  (Marshall et al., 2020), a continuously updated database of RCTs, as potential evidence. We provide experts with these RCT abstracts and their publication dates. We used a dense retrieval system using state-of-the-art embedding models. We provided a more detailed description of this retrieval methodology in Appendix F. We do not evaluate the feasibility of manual evidence search in this study. However, experts do provide judgments on how relevant retrieved abstracts are to the claim in question. For each RCT abstract, we collected annotations determining its relevance to the claim along four dimensions: Population, Intervention, Outcome (PIO), and overall relevance. Claims are contextualized in their original Reddit posts during annotation. Annotators labeled each dimension as (1) Relevant, (2) Somewhat Relevant, or (3) Irrelevant. If an abstract was deemed overall relevant, annotators highlighted the most relevant text span and assessed whether the trial described in the abstract supports or refutes the claim using the four labels: (1) Supports, (2) Partially Supports, (3) Partially Refutes, and (4) Refutes.
@@
    Expert Support: Optional verification incorporating their clinical experience and opinion. This is especially important for outlandish claims that are unlikely to have any trial evidence.
@@
To justify their veracity label, experts write a paragraph-length explanation (see guideline in Appendix M). Annotators may optionally include a medical addendum detailing clinical practices typically used in response to the claim, providing practical context for users.
Our annotation team consists of five clinical experts, one serving as the medical lead. All experts had experience reviewing medical articles and synthesizing them for biomedical research or patient care (annotator recruitment details are in Appendix H). To leverage their expertise effectively, we conducted the study in three rounds, with changes between rounds detailed in Appendix N. Following  Klie et al. (2024), the two co-first authors held group meetings with the experts during each round to discuss disagreements and refine the annotation guidelines. In total, these meetings spanned four hours.
4 Results and Analysis
@@ 4.1 Challenge 1: Connecting Medical Evidence with Claims
We present the agreement on the final round of annotations for five claims annotated by five experts (totaling 25 separate verifications) in Table 5. This round includes 50 abstracts, with five labels per abstract, and two labels per claim. Despite multiple rounds of expert feedback to improve the annotation task, agreement remained low across all fields. A guideline for two annotators on two classes suggests that a Fleiss’ $\kappa$ score of 0.21-0.40 represents fair agreement, 0.41-0.60 moderate agreement, and 0.61-0.80 substantial agreement (Landis and Koch, 1977). Since $\kappa$ scores tend to be higher with fewer categories (Sim and Wright, 2005), we consider a reasonable $\kappa$ score for our task with 4+ labels to be around 0.5.
Most instances—20 out of the 25 expert judgments—were judged to have “No Relevant Abstracts”, indicating that the claims were unverifiable. For three of the five claims in this final round, all experts agreed they were unverifiable (see example in LABEL:table:trikafta_example). This high rate of unverifiable claims underscores a broader challenge that “claims in the wild” are so often unverifiable given even the best conditions: The RedHot claims are annotated to “suggest (explicitly or implicitly) a causal relationship between an Intervention and an Outcome” (Wadhwa et al., 2023); the evidence is retrieved with a state-of-the-art system.
Grapefruit
Post (r/Epilepsy): Grapefruit and siezures
Guys, I am wondering if you have any issues or know about interactions between oxcarbazepine and/or levetiracetam and grapefruit? I believe it may make those medications work differently, but I am not sure.
Population: Epilepsy patients (implied by the subreddit r/Epilepsy and the mention of seizure medications),
Intervention: Grapefruit consumption (in interaction with oxcarbazepine and/or levetiracetam),
Outcome: Medication efficacy (i.e., how the medications work)
Table 3: This claim is unverifiable because no RCTs have examined the interactions between grapefruit, Oxcarbazepine, and epilepsy, and conducting such a study may be infeasible.
Dandruff and Trikafta
Population: Patients with Cystic Fibrosis (implied by the subreddit r/CysticFibrosis),
Intervention: Trikafta (a medication),
Outcome: Flaky scalp or dandruff
@@
Table 5: Inter-annotator agreement for each portion of the fact-checking pipeline. Blue: abstract level labels, pink: synthesis level labels.
@@
These issues highlight the difficulty of collecting evidence that is directly relevant to claims people make (on social media). Medical evidence, especially high-quality evidence, is bounded and restricted by standards for feasibility and ethics in a way that covers all the possible queries from the public is impossible. Prior to annotation, we used an automated method to filter out claims that could not be verified by RCTs (detailed in Appendix K). This issue of unverifiability remained despite this effort. Part of solving the issue lies in addressing this issue of expanding the pool of evidence while ensuring it is still trustworthy (See Section 5.4). Additionally, systems should tackle how to best handle the inevitable case in which a claim is unverifiable. We discuss a guided retrieval approach to this in Section 5.2.
@@
Post (r/ADHD): Hello, menstruating people! How do your cycle and ADHD influence each other and how do you deal with it?
EDIT: After getting your responses I am reflecting again how medicine does not give a shit about women. It’s truly insane. Thank you! Hello! I have never paid too much attention to my menstrual cycle since it was never particularly bothersome. Now that I take methylo I feel big changes in how I function during the cycle. Like last 10 days of the cycle, my medication kind of stops working… That is like 1/3 of the time. I know it’s still better than without meds nevertheless, it makes establishing a routine quite challenging. My doc suggested trying contraceptive pills, but I am not even sexually active ATM so taking more medication, with potential side effects, does not excite me. I know there are herbs that are proven to be helping with regulating the cycle but I don’t know if they would help with ADHD symptoms? Any tips?
Population: People with ADHD, Intervention: Herbs, Outcome: Regulating the menstrual cycle
@@
Post (r/CysticFibrosis): Anyone with sinus issues drinking pineapple juice?
It’s a weird question, but I saw a post about pineapple juice being good for sinus issues (helps with the inflammation) and just wondered if anyone has done this? Some people were commenting about the high sugar content in pineapple juice not being good, but they get around that by taking a supplement instead of drinking the juice. Anyone?
Population: Patients With Cystic Fibrosis, Intervention: Pineapple Juice, Outcome: Reduced Inflammation/Fewer Sinus Issues
@@
Table 6: Examples of underspecified claims.
@@ 5 How Can We Address These Challenges?
Figure 2: This figure illustrates the communication model for fact-checking, where the system engages the patient in dialogue—asking clarifying questions, filling contextual gaps, and verifying the claim and addressing any misconceptions.
To address the challenges in 4, we argue that fact-checking alone is insufficient: A communication model is required to mirror a dialogue between a healthcare provider and a patient. This frames fact-checking as a dialogue aimed at addressing patient’s information needs. A vision of the user interaction with the communication model is shown in 2.
@@ 6 Conclusion
We have presented an analysis of the construct validity of existing end-to-end automatic medical fact-checking systems, with expert engagement in all key aspects of the system. Our work highlights the unique challenges of automated medical fact-checking, showing that it should be approached with user interaction in mind, and not as an end-to-end system. We proposed a communication model to clarify underspecifications and guide unverifiable claims, aiming to improve user outcomes and real-world utility. We hope this work inspires further exploration of human-in-the-loop systems for medical fact-checking.
@@
Due to factors in time and cost, our annotation study is limited in the number of claims experts annotated. However, we were still able to collect a wealth of information from these annotations and from the insightful discussions with our expert annotators. Ultimately, the fundamental issues with the fact-checking process that we present in this paper acted as a barrier in continuing with a more large-scale annotation study. We used an automatic document retrieval system in lieu of a manual effort from annotators to find relevant abstracts in relation the claim. Therefore, it is possible for such a method to fail to identify the most relevant abstracts with respect to the claim. We did this to avoid overwhelming annotators and because the focus of the study is not on the retrieval of medical documents, but on the process of synthesizing such documents to verify a claim. In fact, a diversity of relevance in the abstract collection would have lent more insight into how such evidence is supposed to be synthesized. However, the fundamental issues we detail in this paper overshadowed and impeded in being able to do such an analysis.
Acknowledgments and Disclosure of Funding

### v3

@@ Abstract
Technological progress has led to concrete advancements in tasks that were regarded as challenging, such as automatic fact-checking. Interest in adopting these systems for public health and medicine has grown due to the high-stakes nature of medical decisions and challenges in critically appraising a vast and diverse medical literature. Evidence-based medicine connects to every individual, and yet the nature of it is highly technical, rendering the medical literacy of majority users inadequate to sufficiently navigate the domain. Such problems with medical communication ripens the ground for end-to-end fact-checking agents: check a claim against current medical literature and return with an evidence-backed verdict. And yet, such systems remain largely unused.
In this position paper, developed with expert input, we present the first study examining how clinical experts verify real claims from social media by synthesizing medical evidence. In searching for this upper-bound, we reveal fundamental challenges in end-to-end fact-checking when applied to medicine: Difficulties connecting claims in the wild to scientific evidence in the form of clinical trials; ambiguities in underspecified claims mixed with mismatched intentions; and inherently subjective veracity labels. We argue that fact-checking should be approached and evaluated as an interactive communication problem, rather than an end-to-end process. Our data and code is available at https://github.com/SebaJoe/decide-less-communicate-more.
Decide less, communicate more: On the construct validity of end-to-end fact-checking in medicine
Sebastian Joseph [^1] footnotemark: 1 Lily Chen2thanks: Equal Contribution Barry Wei3 Michael Mackert1 Iain J. Marshall4 Paul Pu Liang2 Ramez Kouzy5 Byron C. Wallace6 Junyi Jessy Li1 1The University of Texas at Austin, 2Massachusetts Institute of Technology 3Indiana University School of Medicine, 4King’s College London 5The University of Texas MD Anderson Cancer Center, 6Northeastern University {sebaj, mackert, jessy}@utexas.edu, {l1ly, ppliang}@mit.edu, barrwei@iu.edu iain.marshall@kcl.ac.uk, rkouzy@mdanderson.org, b.wallace@northeastern.edu
@@ 1 Introduction
Decision making in medicine is personal, intimate, and high-stakes. Traditionally the patient—often a lay person unfamiliar with medicine—converses with their care providers about questions about their health. However, the reality is far from this picture: most Americans resort to the web when they have a health-related question (Fox and Duggan, 2013). Today, social media and AI have made medical knowledge seemingly accessible. But claims made by others on the web (or by a chatbot) can be inaccurate or inapplicable. This, combined with eroding health literacy (Champlin et al., 2017), have led to challenges in public health (Hassan et al., 2015) as well as patient-provider sessions.
@@
Dataset Domains Source Labels Explanations Evidence Type Claim Example
@@
Table 1: Comparative overview of related work in medical and health fact-checking. N/A indicates components not applicable to the task (e.g., RedHOT does not perform claim verification).
This position paper highlights practical gaps in AI-driven fact-checking systems that cannot be addressed by “building a better system” alone. We focus on real-world medical claims from social media, preserving their original context, and contend that fact-checking systems should mirror how experts (e.g., physicians, care providers) evaluate and respond to such claims. As part of our exploration, we designed an annotation study that examines how medical experts verify claims using retrieved medical evidence. Experts were asked to assess medical claims present on social media (i.e., Reddit forums about a particular medical condition) by synthesizing retrieved randomized controlled trial (RCT) abstracts and explaining their judgments. This provides an idealized upper bound for systems verifying “in the wild” claims. However, we highlight fundamental obstacles that challenge the construct validity of end-to-end automated systems for fact-checking: given a claim, provide a veracity judgment. We identify inherent difficulties in this setup for even domain experts, including: connecting claims with evidence; ambiguity from underspecified claims leading to valid yet contradictory interpretations; and challenges in achieving annotator consensus due to the inherent subjectivity of veracity labels. These issues suggest that the existing framing fact-checking of as an end-to-end classification task is inadequate for real-world settings, which may explain in part why such systems have not been put into wide use.
2 Background: medical claim-checking
@@
Figure 1: Overview of our AI-in-the-loop expert study pipeline. Given a claim from a subreddit, we extract PIO elements and automatically retrieve evidence. The evidence and its context are presented to a medical expert, who provides a veracity judgment and grounded rationale.
@@ 3 Case Study Methods
To reveal the gaps in end-to-end fact-checking systems, we set out to establish an “idealized” scenario: given the task construction for automatic fact-checking systems discussed in Section 2, we asked domain experts to perform all subtasks that are delegable to humans. Thus, we formulate the following study for expert medical claim checking using the AI-in-the-loop pipeline shown in Figure 1.
Claim source
@@
We use claims from the Reddit Health Online Talk (RedHOT) corpus (Wadhwa et al., 2023), which contains 22,000 annotated posts from Reddit across 24 health conditions. RedHOT defines a claim as a statement indicating (often only implicitly) a causal relationship between an intervention and an outcome. To help experts contextualize these claims, we provide the full post and Population, Intervention, Outcome (PIO) descriptors. (See Appendix C, Appendix I for derivation details.) Note that we leave off the Comparators because in practice claims rarely mention the comparator explicitly (e.g., “Vitamin C cured my flu”). Given a contextualized medical claim, a medical expert constructs a hierarchy of evidence (Guyatt et al., 2008a, b, 2011) based on relevance and quality, assesses the claim’s veracity, and provides a grounded explanation.
Annotation task setup
@@
The overall task, without any aid, places a high cognitive load on experts who must comb through and synthesize multiple pieces of evidence (Juneja and Mitra, 2022). With the aim of easing this, we automated several steps of the fact-checking process and integrated them into an intuitive web-based annotation interface (see Appendix G). These features enable experts to focus on critical aspects: evaluating evidence relevance and synthesizing it to support or refute claims. Detailed annotation guidelines are provided in Appendix E.
@@
For each claim, we used its PIO elements to automatically retrieve ten published relevant RCT abstracts from Trialstreamer  (Marshall et al., 2020), a continuously updated database of RCTs, as potential evidence. We provide experts with these RCT abstracts and their publication dates. We used a dense retrieval system using state-of-the-art embedding models. We provided a more detailed description of this retrieval methodology in Appendix K. We do not evaluate the feasibility of manual evidence search in this study. However, experts do provide judgments on how relevant retrieved abstracts are to the claim in question. For each RCT abstract, we collected annotations determining its relevance to the claim along four dimensions: Population, Intervention, Outcome (PIO), and overall relevance. Claims are contextualized in their original Reddit posts during annotation. Annotators labeled each dimension as (1) Relevant, (2) Somewhat Relevant, or (3) Irrelevant. If an abstract was deemed overall relevant, annotators highlighted the most relevant text span and assessed whether the trial described in the abstract supports or refutes the claim using the four labels: (1) Supports, (2) Partially Supports, (3) Partially Refutes, and (4) Refutes.
@@
    Expert Support: Optional verification based on clinical expertise, particularly for claims unlikely to be studied in clinical trials.
@@
To justify their veracity label, experts write a paragraph-length explanation (see guideline in Appendix F). Annotators may optionally include a medical addendum detailing clinical practices typically used in response to the claim, providing practical context for users.
Annotation team
Our annotation team consists of five clinical experts, one serving as the medical lead. All experts had experience reviewing medical articles and synthesizing them for biomedical research or patient care (annotator recruitment details are in Appendix H). To leverage their expertise effectively, we conducted the study in three rounds, with changes between rounds detailed in Appendix D. Following  Klie et al. (2024), the two co-first authors held group meetings with the experts during each round to discuss disagreements and refine the annotation guidelines. In total, these meetings spanned four hours.
Across the three rounds, each of the five experts annotated 20 unique claims, yielding 1,000 abstract-level annotation instances (10 abstracts per claim) and 100 synthesis-level explanations.
4 Challenges for end-to-end fact-checking
In this section, we present the fundamental challenges revealed by our expert annotations and inputs, gathered over multiple rounds of discussion.
@@ 4.1 Challenge 1: Connecting Medical Evidence with Claims
We present the inter-annotator agreement on the final round of annotations for five claims annotated by five experts (totaling 25 separate verifications) in Table 2. This round includes 50 abstracts per expert (250 total), with five labels per abstract, and two labels per claim. Despite multiple rounds of expert feedback to improve the annotation task, agreement remained low across all fields. A guideline for two annotators on two classes suggests that a Fleiss’ $\kappa$ score of 0.21-0.40 represents fair agreement, 0.41-0.60 moderate agreement, and 0.61-0.80 substantial agreement (Landis and Koch, 1977). Since $\kappa$ scores tend to be higher with fewer categories (Sim and Wright, 2005), we consider a reasonable $\kappa$ score for a task with 4+ labels to be approximately 0.5.
@@
Table 2: Inter-annotator agreement for each portion of the fact-checking pipeline. Blue: abstract-level labels; pink: synthesis-level labels.
Most instances—20 out of the 25 expert judgments—were labeled “No Relevant Abstracts”, indicating that the claims were unverifiable. For three of the five claims in this final round, all experts independently labeled them “No Relevant Abstracts” (see Section A for an illustrative example). This high rate of unverifiable claims underscores a broader challenge: even when claims are annotated to suggest a causal relationship between an Intervention and an Outcome (Wadhwa et al., 2023) and are paired with state-of-the-art evidence retrieval, they are often unverifiable in practice.
Grapefruit
Post (r/Epilepsy): Grapefruit and siezures Guys, I am wondering if you have any issues or know about interactions between oxcarbazepine and/or levetiracetam and grapefruit? I believe it may make those medications work differently, but I am not sure.
Population: Epilepsy patients (implied by the subreddit r/Epilepsy and the mention of seizure medications)
Intervention: Grapefruit consumption (in interaction with oxcarbazepine and/or levetiracetam)
Outcome: Medication efficacy (i.e., how the medications work)
Table 3: An example of an unverifiable claim, as no RCTs have examined interactions between grapefruit, oxcarbazepine, and epilepsy, and such a study may be infeasible.
@@
These issues highlight the difficulty of collecting evidence that is directly relevant to claims people make (on social media). Medical evidence, especially high-quality evidence, is bounded and restricted by standards for feasibility and ethics in a way that covers all the possible queries from the public is impossible. Prior to annotation, we used an automated method to filter out claims that could not be verified by RCTs (detailed in Appendix J). This issue of unverifiability remained despite this effort. Part of addressing this challenge lies in expanding the pool of evidence while ensuring it remains trustworthy (see Section 5.4). Additionally, systems should tackle how to best handle the inevitable case in which a claim is unverifiable. We discuss a guided retrieval approach to this in Section 5.2.
@@
Post (r/ADHD): Hello, menstruating people! How do your cycle and ADHD influence each other and how do you deal with it? EDIT: After getting your responses I am reflecting again how medicine does not give a shit about women. It’s truly insane. Thank you! Hello! I have never paid too much attention to my menstrual cycle since it was never particularly bothersome. Now that I take methylo I feel big changes in how I function during the cycle. Like last 10 days of the cycle, my medication kind of stops working… That is like 1/3 of the time. I know it’s still better than without meds nevertheless, it makes establishing a routine quite challenging. My doc suggested trying contraceptive pills, but I am not even sexually active ATM so taking more medication, with potential side effects, does not excite me. I know there are herbs that are proven to be helping with regulating the cycle but I don’t know if they would help with ADHD symptoms? Any tips?
Population: People with ADHD Intervention: Herbs Outcome: Regulating the menstrual cycle
@@
Post (r/CysticFibrosis): Anyone with sinus issues drinking pineapple juice? It’s a weird question, but I saw a post about pineapple juice being good for sinus issues (helps with the inflammation) and just wondered if anyone has done this? Some people were commenting about the high sugar content in pineapple juice not being good, but they get around that by taking a supplement instead of drinking the juice. Anyone?
Population: Patients With Cystic Fibrosis Intervention: Pineapple Juice Outcome: Reduced Inflammation/Fewer Sinus Issues
@@
Table 4: Examples of underspecified claims and corresponding expert feedback.
@@ 5 How Can We Address These Challenges?
Figure 2: This figure illustrates the communication model for fact-checking, where the system engages the patient by asking clarifying questions, filling contextual gaps, and verifying claims while addressing misconceptions.
To address the challenges in Section 4, we argue that fact-checking alone is insufficient: A communication model is required to mirror a dialogue between a healthcare provider and a patient. This frames fact-checking as a dialogue aimed at addressing patient’s information needs. A vision of the user interaction with the communication model is shown in Figure 2.
@@ 6 Conclusion
In this position paper, we present an analysis of the construct validity of existing end-to-end automatic medical fact-checking systems, with expert engagement in all key aspects of the system. Our work highlights the unique challenges of automated medical fact-checking, showing that it should be approached with user interaction in mind, and not as an end-to-end system. We propose a communication model to clarify underspecifications and guide unverifiable claims, aiming to improve user outcomes and real-world utility. We hope this work inspires further exploration of human-in-the-loop systems for medical fact-checking.
@@
Due to time and cost constraints, our annotation study includes a limited number of claims. However, each claim was evaluated against 10 retrieved abstracts, resulting in a substantial volume of evidence-based annotations. These annotations and expert discussions revealed fundamental challenges in medical fact-checking that make large-scale annotation difficult to define in a principled way.
We used an automatic document retrieval system instead of manual search by annotators to identify relevant abstracts for each claim. While this approach may fail to retrieve the most relevant evidence, retrieval performance was not the focus of this study. Rather, our goal was to examine how retrieved medical evidence is synthesized to verify claims, while avoiding undue burden on expert annotators.
These limitations reflect broader challenges in defining and evaluating medical fact-checking tasks in practice.
Ethical Considerations
The posts found within the RedHOT dataset contain health-related comments that are inherently sensitive. To respect this sensitivity, the authors of the RedHOT dataset notified all users of their inclusion in this dataset and provided them with the opportunity to opt-out. They also did not release the data directly, but instead provided a script to download content from Reddit so that individuals can remove their post in the future. In this work, we directly release a small subset of these posts from RedHOT that we used in our annotation study. In our released data, we do not reveal the username of the author of the post. We only include the text from the post and information about the subreddit in which it was found. Considering the measures taken by the authors of the RedHOT dataset and the fact that these posts have been publicly available on Reddit for at least more than a year, we believe it is safe to publicly release this data.
We have consulted with our Institutional Review Board (IRB) about the nature of our work and confirmed that the use of the RedHOT data and the subsequent annotation study using these data do not constitute research of human subjects. However, we acknowledge that certain uses of this data may be considered sensitive. We strongly encourage researchers to obtain prior approval from their own IRB regarding the intended use of the data released by this work.
We also consider ethical aspects of the annotation process. Annotators consented to participate and were compensated at a competitive hourly rate (see Appendix H for details).
Acknowledgments